• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name M. Stevens

    Log entry time 08:06:56 on March 19, 2007

    Entry number 196564

    keyword=TECH Left Q1 trip, odd reset

    Left q1 arbitrarily tripped. No disturbances in the cryogenics. Failure mode: Magnet appeared to be communicating; in "off" mode we could change the polarity but we couldn't ramp the magnet; Unresetable "Spare Interlock" and "Max Current Set" flags. Reset everything on the gui and the green button in the counting house center room to no avail. On a long shot we reset IOCHA48, and that worked, but we still see the above flags. These flags are informational only and not among the hardware interlock faults. We ramped the magnet to the needed current and resumed operations.

    A copy of this log entry has been emailed to: spiegel@jlab.org, fansler@jlab.org, folts@jlab.org, stevensm@jlab.org